OpenStudy (purple_pink):

In one basketball season, Tim made 75% of the free throws he attempted. Tim made 36 free throws. How many free throws did he attempt? A. 27 B. 48 C. 61 D. 75

OpenStudy (purple_pink):

@Zelda @rebeccaxhawaii

OpenStudy (anonymous):

If he made 75% of his free throws. And he did 36 of them, it's 36/X So you do: \[\frac{ 3 }{ 4 } = \frac{ 36 }{ x }\] Solve for x.

OpenStudy (purple_pink):

thanks! :) @Zelda

OpenStudy (purple_pink):

x=48